### v3

# Symbolic Mixture-of-Experts: Adaptive Skill-based Routing for Heterogeneous Reasoning

Abstract
Combining existing pre-trained expert LLMs is a promising avenue for scalably tackling large-scale and diverse tasks. However, selecting experts at the task level is often too coarse-grained, as heterogeneous tasks may require different expertise for each instance. To enable adaptive instance-level mixing of pre-trained LLM experts, we propose Symbolic-MoE, a symbolic, text-based, and gradient-free Mixture-of-Experts framework. Symbolic-MoE takes a fine-grained approach to selection by emphasizing skills, i.e., specialized subcategories such as algebra in mathematics. We propose a skill-based recruiting strategy that dynamically selects the most relevant set of expert LLMs for diverse reasoning tasks based on their strengths. Each selected expert then generates its own reasoning, resulting in $k$ outputs from $k$ experts, which are then synthesized into a final high-quality response by an aggregator, which is chosen based on its ability to integrate diverse outputs. We show that instance-level expert selection improves performance by a large margin but – when implemented naively – can introduce a high computational overhead due to the need for constant model loading and offloading. To address this, we implement a batch inference strategy that groups instances based on their assigned experts, ensuring each model will only be loaded once. This allows us to integrate 16 models on a single GPU with a time cost comparable to prior multi-agent baselines using 4 GPUs. Through extensive evaluations on diverse benchmarks (MMLU-Pro, GPQA, AIME, and MedMCQA), we show that Symbolic-MoE outperforms prior multi-agent approaches, with an absolute average improvement of $8.15\%$ over the best baseline. Moreover, Symbolic-MoE generalizes well to unseen tasks and removes the need for expensive multi-round discussions, outperforming discussion baselines with less computation. [^1] Project page: https://symbolic-moe.github.io
1 Introduction
A core strength of humans is our ability to communicate and coordinate with each other using language (Clark, 1996; Yow & Lim, 2019; Xu et al., 2023). This allows diverse experts to contribute specialized knowledge towards solving a problem. Like humans, large language models (LLMs) often have differing skills and strengths, derived from differences in their architectures and training regimens. For instance, math-specific models like MetaMath (Yu et al., 2023) or QwenMath (Yang et al., 2024) are post-trained with mathematical reasoning data, making them particularly adept at math tasks – often at the cost of performance on out-of-distribution tasks (Kumar et al., 2022; Chu et al., 2025) like commonsense or medical reasoning (Lobo et al., 2024). Even within specialized domains, differences in pre-training data can lead to nuanced variations in expertise: one math-focused model may excel at algebra, while another is better suited for geometry. This motivates our development of an automated, skill-based framework designed to identify and select the most suitable set of expert models for each problem.
Indeed, combining multiple “expert” models via Mixture-of-Experts (MoE) is well-studied (Jacobs et al., 1991; Eigen et al., 2013) and has been applied widely for large pre-trained models, enabling better performance at a lower computational cost (Shazeer et al., 2017a; Fedus et al., 2022; Riquelme et al., 2021). However, in the conventional MoE settings, experts are typically sub-models, i.e. subsets of parameters within a larger model, where at test time, they are combined in the model’s parameter space. This generally requires end-to-end training from scratch, which is often computationally expensive and precludes the re-use of the vast pool of existing already-trained LLMs. Building on recent efforts in combining a fixed set of models through multi-agent discussions (Chen et al., 2024c; Du et al., 2023; Liang et al., 2023; Wang et al., 2024a), we propose exploring a new training-free paradigm for large-scale MoEs: a symbolic mixture of experts (Symbolic-MoE). Rather than using information encoded in the model’s hidden state, Symbolic-MoE uses symbolic structures in two ways: First, Symbolic-MoE infers a set of discrete skills needed to solve a problem, measuring the abilities of each model in a pool of candidate expert models. It then uses skill-based performance as a “router” to recruit a sparse subset of experts for each problem. Secondly, Symbolic-MoE combines pre-trained experts through a symbolic channel, i.e., language, which is a common protocol already shared by all LLMs. To take advantage of the diverse set of expert LLMs, we must address two key: (1) Effective Expert Selection: Given a large set of existing LLMs, how can we choose the best experts for each instance? (2) Scalable Expert Mixing: How can we serve a large number of experts (e.g. $16$) without increasing the demand for GPUs?
(1) Effective Expert Selection: The increasing diversity of benchmarks (Miranda et al., 2024) and the growing number of models means that experts must be selected not at the level of tasks, but at the level of individual queries. Even at the task level, manual selection can be labor-intensive, and the performance of multi-agent frameworks can be sensitive to the agents chosen for a task (Chen et al., 2024c; Liang et al., 2023; Wang et al., 2024b). For instance, as shown in Fig. 1 (a), while a given subset of models may perform well on math tasks on average, their proficiency in specific subfields like algebra or probability might vary – that is, using a fixed subset of models on all math samples might hurt performance on particular subtasks. This underscores the need for an automated, fine-grained selection mechanism, as shown in Fig. 1 (b).
Figure 1: (a) In prior work, a fixed set of task-level experts is recruited to solve mathematical problems, while heterogeneous questions may differ in the skills required to solve them (e.g., Q1 requires algebra, while Q2 focuses on probability). The recruited experts generate outputs for multiple rounds, making these methods inefficient. (b) In contrast, Symbolic-MoE adaptively recruits instance-level experts based on skills needed (“Algebra” experts for Q1 and a different set of “Probability” experts for Q2). By generating only a single round of responses with an aggregator to synthesize the final output, our approach is both more performant and more efficient.
(2) Scalable Expert Mixing: Past work has often relied on multiple rounds of inference, leading to significant GPU demands. Moreover, it does not scale to a dynamic setting like the one we consider, where the number of GPUs required would be equal to the number of potential models available (in our case, $16$), making this option prohibitively expensive. We instead propose a batch inference mechanism (see Fig. 3 (III)). Specifically, the samples are grouped into batches per model. We then run all queries assigned to a given model in a single batch, which is far faster than sequential processing. While this strategy accommodates up to $16$ models on a single GPU, it can also be parallelized across multiple GPUs. This flexibility ensures both speedups with increased computing power, and accessibility for users with limited hardware resources.
We evaluate Symbolic-MoE on diverse benchmarks, including MMLU-Pro (Wang et al., 2024c), GPQA (Rein et al., 2023), AIME 2024 (MAA, 2024), and MedMCQA Pal et al. (2022), using a diverse model pool. We show that our automated skill-based recruiting yields an average accuracy improvement of $8.15\%$ over the best multi-agent baseline. Moreover, despite primarily using LLMs with 7-8 billion (B) parameters, Symbolic-MoE achieves comparable performance with 70B models, and on average, outperforms strong proprietary models like GPT4o-mini (OpenAI, 2024), Gemini 1.5 Pro (Team et al., 2024a), and DeepSeek-V3 (DeepSeek-AI et al., 2025b). Also, Symbolic-MoE consistently surpasses all baselines, whereas the strongest baseline changes across tasks. Thus, Symbolic-MoE eliminates the need for the user to evaluate and compare a large number of possible baselines for each task. Notably, using a single GPU, Symbolic-MoE has $44\%$ less inference run-time than a mixture-of-agents baseline (Wang et al., 2024a); when four GPUs are available for both methods, we obtain an almost 2×2$\times$2 × speedup. Finally, our analysis shows that Symbolic-MoE generalizes well to unseen tasks, and selecting a task-specific aggregator achieves performance comparable to multi-round discussion while requiring substantially less compute.
2 Related Work
Mixture-of-Agents. Traditional Mixture-of-Experts (MoE) models (Jacobs et al., 1991; Jordan & Jacobs, 1994; Chen et al., 1999; Yuksel et al., 2012) distribute computation across multiple “expert” submodels, with growing interest in sparsity-driven approaches. The Sparse MoE (SMoE) approach (Shazeer et al., 2017a) improves efficiency by activating only the most relevant experts per input, enhancing scalability for high-dimensional data. This method has been widely applied in vision tasks (Riquelme et al., 2021; Wang et al., 2020; Yang et al., 2019; Abbas & Andreopoulos, 2020), language tasks (Lepikhin et al., 2021; Zhang et al., 2021; Zuo et al., 2022; Jiang et al., 2021) and multimodal learning (Kudugunta et al., 2021; Yun et al., 2024). Unlike Symbolic-MoE, standard MoE approaches require experts to be trained jointly, with communication taking place in the parameter spaces of the submodels. On the test-time mixture side, MoA (Wang et al., 2024a) combine LLM agents into ensembles that rely on a fixed set of agents. This approach requires multiple rounds of generation and aggregation before producing a final answer. Similarly, Self-MoA  (Li et al., 2025) suggest that optimal performance can be achieved by invoking the task-best model multiple times alongside the task-best aggregator. Our work differs from MoA and Self-MoA by introducing adaptive, instance-level, skill-based routing while avoiding costly multi-model discussions in favor of streamlined aggregation. We also find that mixing different LLMs is advantageous when paired with effective routing and aggregation strategies. Moreover, we show that the best aggregator for a task is not necessarily the best-performing model overall, but can be identified through a synthetic task we introduce, designed to evaluate aggregation effectiveness.
Multi-Agent Reasoning. Multi-agent reasoning has emerged as a promising paradigm for enhancing complex problem-solving and decision-making in AI systems. Early approaches employed reinforcement learning-based coordination (Lowe et al., 2017; Foerster et al., 2018; Jaques et al., 2019), while recent efforts leverage LLM-based multi-agent frameworks. One line of research explores student-teacher paradigms (Magister et al., 2022; Fu et al., 2023; Ho et al., 2022; Du et al., 2023; Chen et al., 2024a), where reasoning capabilities are distilled from stronger to weaker agents. Another approach investigates multi-agent debate frameworks, where agents interact to refine arguments and enhance collective decision-making; this has been explored with multiple instances of a single model (Liang et al., 2023; Xiong et al., 2023; Chan et al., 2023) or debates between multiple LLM types (Chen et al., 2024c). In both cases, the set of models is predefined by the user. In contrast, our approach automatically selects models based on inferred skills. Additionally, our framework achieves superior efficiency by avoiding multi-round discussions while still outperforming debate-based methods.
Figure 2: Overview of Symbolic-MoE. (a) Preprocessing: Given a validation set and a pool of agents, we create model profiles and select an aggregator. (b) Inference-Time: For each test example, Symbolic-MoE activates the most relevant models (experts) based on skill-based routing, using model profiles determined during preprocessing. These models generate CoT responses, which the aggregator (chosen based on its ability to select correct answers) synthesizes into a final answer.
3 Methodology
3.1 Problem Setup
Given a pool of $n$ models =$\mathcal{M}=\{M_{i}\}_{i=1}^{n}$, where each model represents a distinct LLM with potentially different pre-training datasets and architectures, our goal is to optimize performance through dynamic allocation – solving each problem with the most suitable subset of $k$ models, allowing experts to combine information to enhance reasoning. To achieve this, we assume access to a small validation set to obtain (1) model profiles $P_{i}\;\forall i\leq n$, and (2) aggregator profiles that benchmark the ability of each model to act as an aggregator. We use these profiles to recruit experts (at the instance level) and to select the aggregator (at the task level).
3.2 Preprocessing
3.2.1 Model Profile Creation
To recruit the $k$ most suitable experts for a given question, we assess each model’s specialized skills across various problem-solving domains, illustrated in Fig. 2 (a). This is done by evaluating their performance on the validation set for each task (see Table 13 for sizes), thereby constructing a model profile $P_{i}$ for each model $M_{i}$. For each question in the validation set, we first prompt an LLM – referred to as the “Keyword LLM” – to identify the essential skills required to solve the problem. For consistency, we use Qwen2.5-7B-Instruct (Qwen Team, 2024) as the Keyword LLM. To reduce noise, we generate keyword annotations for each question five times, and retain only those that appear more than once across the whole validation set. These extracted skills represent core knowledge areas necessary for solving the problem – for instance, a given college-level math problem may require skills such as algebra, calculus, and geometry. Once all questions are annotated with their required skills, each model $M_{i}$ in the pool attempts to solve them using Chain-of-Thought reasoning (Wei et al., 2022). A correct answer increases the score of each associated skill by $+1$, while an incorrect answer results in a $-1$ penalty. At the end of this process, each model has a profile $P_{i}$ represented as a dictionary, e.g., {‘Algebra’: 10, ‘Biology’: 3, ‘Chemistry’: -6, …}.
3.2.2 Aggregator Selection
An aggregator is a model that consolidates $k$ outputs into a single high-quality response. Our pilot experiments, along with findings from Wang et al. (2024a) and Li et al. (2025), indicate that the aggregator model plays a crucial role in the final performance, and selecting the most effective model for aggregation is non-trivial. We find that a strong reasoning model is not necessarily a strong aggregator and vice versa; qualitatively, we show this later in Table 6. We also find that adaptively selecting an aggregator on the instance level based on model profiles is less effective, motivating us to choose the aggregator based on its ability to select correct answers. To identify the best aggregator per task, we construct a synthetic task using the same validation set. From the profile creation process, we obtain outputs from all models, some correct and some incorrect. For each question, we sample one correct reasoning chain and two incorrect ones, structuring the input as follows: {question}, {correct_CoT}, {incorrect_CoT}, {incorrect_CoT}. We shuffle the order of the CoTs and instruct each model to act as an aggregator (using the prompt shown in Appendix L), synthesizing a final output with a predicted answer. We then benchmark each model’s aggregation ability and select the best-performing aggregator for each task.
3.3 Inference
3.3.1 Skill-based Recruiting
At inference time (see Fig. 2 (b)), we follow the same keyword annotation procedure as in Section 3.2.1 to generate relevant keywords for the test sample. To align a test sample’s keywords with those in the model profiles, we employ Sentence-BERT (Reimers & Gurevych, 2020) to match keywords via the cosine similarity between their embeddings. Next, expert recruitment is performed by selecting the top $k$ models whose profiles best match the required skills of the test sample. This is determined by two factors: (1) local suitability score and (2) global competency. For each model $M_{i}$, its local suitability score for a test sample $q$, $\mathcal{S}(M_{i},q)$ is computed as the sum of its skill scores over the set of keywords needed for $q$ (denoted as $K_{q}$). It can be expressed as follows:
$\mathcal{S}(M_{i},q)=\sum_{k_{j}\in K_{q}}s^{(i)}_{k_{j}}$
where $s^{(i)}_{k_{j}}$ represents the score of model $M_{i}$ for the $j$-th skill in the test sample $q$. This results in an model ranking distribution $\mathcal{D}_{q}$ for each test sample $q$: $\mathcal{D}_{q}=\{\mathcal{S}(M_{1},q),\mathcal{S}(M_{2},q),...,\mathcal{S}(M_{n},q)\}$.
Intuitively, suppose $M_{1}$ has scores of $+3$, $+5$, and $-2$ for algebra, calculus, and geometry, respectively, which are needed for a given sample; its total score for this sample would be $3+5-2=6$. Calculating this score for all models yields a distribution of model strengths for the given sample, e.g., {$M_{1}$: 6, $M_{2}$: 3, …, $M_{n}$: -10}, which is the ranking of how suitable a model is for a sample.
To account for each model’s overall strength in a task, i.e., global competency, we compute each model’s total score across all keywords in their profile, and normalize it by the total sum of all agents’ scores. We denote this global strength as $\gamma_{i}$, representing a model’s overall task performance relative to others. Finally, the expert selection is performed by sampling from the product of the local suitability score, $\mathcal{S}(M_{i},q)$ (from a model rank distribution $\mathcal{D}_{q}$) and the global competency $\gamma_{i}$. That is, the relevance score of a model $M_{i}$ for a test sample $q$ is: $w^{(i)}_{q}=\gamma_{i}\mathcal{S}(M_{i},q)$. We apply a softmax function with the temperature set to $0.5$ to this distribution $\{w^{(i)}_{q}\}_{i=1}^{n}$, and then sample $k$ experts for each problem, i.e.,
$E^{(i)}_{q}\sim\text{Categorical}(w^{(1)}_{q},w^{(2)}_{q},...,w^{(n)}_{q}),~{}i=\{1,2,...,k\}$
To enhance efficiency, we trim low-frequency experts, i.e., those who appear in fewer than $5\%$ of test cases. For example, given a test set with 100 samples, where 3 experts are recruited per sample (totaling 300 selections), any expert appearing fewer than $300\times 5\%=6$ times is discarded and resampled from the remaining higher-frequency experts. To visualize this, we provide the expert distribution before and after trimming in Fig. 4.
3.3.2 Final Answer Generation
[…]
3.3.3 Scalable Batched Inference
In our experiments, we mostly consider 7B–8B parameter LLMs, which have a substantial memory footprint. Due to the adaptive nature of the recruitment process, the set of participating LLMs may change dynamically for different problems. For instance, one sample may require Qwen, Llama, and Mistral, while another may need Gemma, Exaone, and Phi. A naive implementation of this approach can lead to high latency, particularly when the required models change frequently. In such cases, the computational cost includes not only inference but also data transfers across GPUs. To mitigate these computational challenges, Symbolic-MoE introduces a novel batching strategy to maximize throughput. Specifically, for a given set of instances, we precompute (using inferred skills) which LLMs will be called for each instance. We then group instances based on their required experts, as illustrated in Fig. 3 (III) and Algorithm 1 in Algorithm 1. In other words, each active expert receives all relevant instances at once, ensuring that each expert is loaded only once per batch. This enables efficient batched inference on a single GPU while supporting a pool of $16$ LLMs. Moreover, this approach is flexible, as more GPUs can further accelerate inference through parallelization.
4 Results and Analysis
4.1 Experimental Setup
We consider $16$ LLMs ranging from 3.5B to 12B parameters, with most models falling in the 7–8B range. These include general-purpose instruction-tuned models, domain-specific fine-tuned variants on math and biology, and models distilled from DeepSeek R1’s trajectories (DeepSeek-AI et al., 2025a). A full list of models is provided in Table 9. We measure performance on a diverse range of datasets, chosen to require expertise in a number of domains. First, we consider MMLU-Pro (Wang et al., 2024c), which is a harder version of MMLU (Hendrycks et al., 2021a), containing a variety of questions across 14 college-level subjects. Given its large test set of 12,000 samples and the computational cost of evaluating proprietary models, we employ stratified sampling to create a subset of 2,100 samples, ensuring each category contains 150 samples. We also evaluate on AIME 2024, which is a challenging mathematics competition dataset containing math Olympiad problems. For more science-specific reasoning, we test on GPQA Diamond (Rein et al., 2023), which contains questions across a variety of science fields written by experts, explicitly designed to be difficult to answer even by skilled humans with web access. Finally, we include MedMCQA (Pal et al., 2022), which covers questions sourced from medical entrance exams across 21 medical subjects. For each dataset, we sample around 350 samples as the validation set to create the model profiles. [^2] For AIME, we sample validation questions from prior years’ problems (2012-2023). Full dataset statistics are provided in Table 13, and the model pool we consider is shown in Appendix D. For single-model baselines, we use the strongest task-specific LLM, while for multi-model baselines, we select the top three models per task. These selections, like those in Symbolic-MoE, are made based on validation performance, which is summarized in Table 10. More details on baselines and their implementations can be found in Appendix B and Appendix C.
Category Method Model MMLU-Pro AIME GPQA MedMCQA Avg.
Zero-Shot CoT GPT4o-mini 63.95 10.00 42.93 68.18 46.27
Zero-Shot CoT Gemini 1.5 Pro 76.38 36.67 61.62 72.68 61.84
Close-Source Single Model Zero-Shot CoT DeepSeekV3 76.29 26.00 60.10 74.09 59.12
Open-Source Zero-Shot CoT Qwen2.5 72B 71.54 $\pm$ 0.88 25.55 $\pm$ 3.85 51.02 $\pm$ 0.27 69.02 $\pm$ 0.32 54.28
70B Model Zero-Shot CoT Llama3.3 70B 69.26 $\pm$ 0.47 32.22 $\pm$ 3.85 51.44 $\pm$ 0.62 59.78 $\pm$ 0.74 53.18
Open-Source 7B Model Zero-Shot CoT QwenR1 52.57 $\pm$ 0.45 55.93 $\pm$ 5.16 44.95 $\pm$ 1.49 38.72 $\pm$ 0.44 48.04
Zero-Shot CoT Task-Best 54.89 $\pm$ 0.53 55.93 $\pm$ 5.16 48.43 $\pm$ 3.10 55.44 $\pm$ 0.50 53.62
Advanced Single Model Self-Refine (SR) Task-Best 53.74 $\pm$ 0.20 53.33 $\pm$ 3.34 50.84 $\pm$ 3.65 49.57 $\pm$ 0.59 51.87
Self-Consistency (SC) Task-Best x5 56.71 $\pm$ 0.14 67.78 $\pm$ 1.57 53.54 $\pm$ 0.36 56.85 $\pm$ 0.11 58.72
Single-Model Multi-Agent Debate Task-Best x3 56.21 $\pm$ 0.55 56.67 $\pm$ 6.67 50.51 $\pm$ 0.51 51.63 $\pm$ 0.80 53.76
Self-MoA Task-Best x3 55.43 $\pm$ 0.72 55.56 $\pm$ 5.09 52.86 $\pm$ 1.46 53.27 $\pm$ 0.60 54.28
Multi-Model Multi-Agent MoA Top-3 61.78 $\pm$ 0.25 41.11 $\pm$ 5.09 52.86 $\pm$ 3.37 59.29 $\pm$ 0.32 53.76
ReConcile Top-3 56.46 $\pm$ 0.10 50.00 $\pm$ 7.20 47.98 $\pm$ 2.32 60.74 $\pm$ 0.43 53.80
Symbolic-MoE Adaptive 63.71 $\pm$ 0.43 68.88 $\pm$ 5.08 57.78 $\pm$ 2.09 59.35 $\pm$ 0.14 62.43
Table 1: Comparison of Symbolic-MoE with single-model and multi-model baselines. Symbolic-MoE outperforms all multi-agent baselines and achieves performance comparable to strong proprietary models like GPT4o-mini, as well as 70B models, while primarily operating with 7-8B models. Notably, no single baseline consistently secures the second-best performance, even when the strongest models for each task are known. In contrast, our method demonstrates robustness, yielding superior results through adaptive expert selection. We bold the best results and underline the second-best (excluding methods using bigger or proprietary models, shown in gray).
4.2 Main Results
We present the main results in Table 1, and summarize the key findings below.
Symbolic-MoE consistently outperforms all baselines. Across all domains, Symbolic-MoE shows superior performance compared to all baselines, beating single-model baselines (e.g., SR, SC) using the best overall model, multi-agent debate with a single model (e.g., Debate, Self-MoA), as well as multi-model multi-agent baselines (e.g., MoA, ReConcile). Symbolic-MoE outperforms the most competitive multi-agent baseline, Self-MoA, by $8.15\%$ (absolute) on average, with consistent improvements across domains (e.g., $8.28\%$ gain on MMLU-Pro, $13.45\%$ on AIME, $4.92\%$ on GPQA, $6.08\%$ on MedMCQA). These gains are also seen when comparing to multi-model baselines like MoA and ReConcile that use the top three strongest models per domain. Symbolic-MoE also substantially outperforms test-time scaling methods, such as SR (Madaan et al., 2023b) and SC (Wang et al., 2023a). Surprisingly, with the task-best model, SC beats multi-agent debate baselines (e.g., Self-MoA, MoA), though it still underperforms Symbolic-MoE by an average of $3.71\%$. This indicates that scaling test-time compute with the task-best model is a simple yet effective way to improve performance, and adaptively selecting suitable experts leads to further improvements.
Symbolic-MoE generalizes well across tasks. No single baseline in Table 1 is universally effective across all tasks. For instance, while MoA performs well on MMLU-Pro, it struggles on AIME; ReConcile excels in MedMCQA but fails to generalize to GPQA. Knowing which method works best for a task is therefore nontrivial, requiring running every method on validation sets to choose from the many settings available. In contrast, Symbolic-MoE consistently delivers strong performance across all domains. Symbolic-MoE especially excels on AIME and GPQA, where SC is a surprisingly strong baseline, and where other strong methods like MoA and Self-MoA fall far behind Symbolic-MoE. Moreover, we see that, while SC with the top model is the most competitive setting on AIME and GPQA, it lags behind other baselines on MMLU-Pro and MedMCQA, where multi-agent baselines perform better. This discrepancy may stem from the broader subject diversity in MMLU-Pro and MedMCQA, where agent collaboration facilitates better consensus, whereas AIME is more math-focused, and favors individual model performance – the task-best model, QwenR1 7B, delivers strong solo performance already. In light of that, we include QwenR1 7B in Table 1, which is a powerful model distilled from DeepSeek-R1’s trajectories (DeepSeek-AI et al., 2025a). While QwenR1 demonstrates exceptional math and code reasoning capabilities ($55.93\%$ on AIME), leading to strong Self-Consistency performance ($67.78\%$), it struggles to generalize to other domains such as MedMCQA. This further underscores the need for a robust and flexible framework like Symbolic-MoE to achieve broad generalization across diverse tasks.
Symbolic-MoE matches strong proprietary models and larger 70B models. In Table 1, we also find that Symbolic-MoE achieves a similar average performance to models that have substantially more parameters. For example, Symbolic-MoE outperforms Llama3.3 70B on AIME and GPQA and roughly matches it on MedMCQA, despite requiring only four 7-8B models (three for the experts and one for the aggregator). Similarly, Symbolic-MoE outperforms or matches a number of strong proprietary models on average – for instance, it matches Gemini 1.5 Pro and outperforms GPT4o-mini, driven in part by sizable gains on AIME and GPQA. These results underscore that, by drawing from a large pool of experts, Symbolic-MoE enables a smaller set of models to achieve performance comparable to models with significantly more parameters, especially when considering a heterogeneous set of tasks like ours.
4.3 Efficiency Anaylsis
Method # GPUs Run Time (s)
Sequential Inference 1 196.92
MoA 1 45.98
MoA 4 21.66
Symbolic-MoE 1 25.76
Symbolic-MoE 4 10.85
Table 2: Efficiency comparison of MoA and Symbolic-MoE on the GPQA test set. Run time is averaged per sample.
Run time efficiency of Symbolic-MoE. We compare Symbolic-MoE’s efficiency to a naive implementation of sequential inference and to that of MoA and in Table 2. We measure the average run-time at inference on GPQA. Unsurprisingly, the sequential inference baseline shows the highest latency, since the model is constantly changing, requiring loading and offloading it for each instance. We also find that Symbolic-MoE operates in $44\%$ less time on a single GPU than MoA, which requires multi-round discussion, with better accuracy. Moreover, Symbolic-MoE running on a single GPU shows a similar time with MoA on 4 GPUs, and furthermore, when we run Symbolic-MoE on 4 GPUs, it results in almost 2×2$\times$2 × speedup over MoA. We provide another efficiency measurement, token count comparison, in Appendix H, where we find that Symbolic-MoE has the best trade-off between token usage and performance among multi-agent baselines.
Discuss Aggr. MMLU-Pro GPQA
Adaptive 59.07 57.01
Adaptive 57.12 58.01
Task-best 57.81 57.78
Task-best 56.67 57.01
Task-specific 63.83 57.72
Task-specific 63.71 57.78
Table 3: Comparison of with and without discussion across varying aggregators. Discussion stabilizes performance with suboptimal aggregators but has little effect with an optimal aggregator.
Methodological efficiency of Symbolic-MoE. Like multi-agent discussion baselines, Symbolic-MoE can also operate in a discussion-based manner. Instead of aggregating initial responses, models first engage in a round discussion before submitting final answers to an aggregator. Table 3 evaluates this approach, comparing the adaptive aggregator (suboptimal) and task-best aggregator (suboptimal), as well as the task-specific aggregator (optimal) on MMLU-Pro and GPQA. Given the optimal aggregator, we see that discussion yields marginal gains on MMLU-Pro ($63.83$ vs. $63.71$) and a modest drop on GPQA ($57.72$ vs. $57.78$). We find that while round-wise discussion does improve performance incrementally, but the final outcome is ultimately determined by the strength of the aggregator. As a result, Symbolic-MoE has less run time than the baselines while maintaining or even surpassing their performance, as demonstrated in Table 2.
4.4 Additional Analysis
Recruiting Strategy Acc.
Top-3 Experts 52.86
Top-5 Experts 47.68
3 Random Experts 42.61
5 Random Experts 44.92
Model Profile (Ours) 57.78
Table 4: Comparison of different recruiting strategies on GPQA.
Utility of the Model Profile. Symbolic-MoE profiles models based on their skills and leverages this information to recruit experts effectively. To underscore the importance of this step, we compare several alternative selection strategies in Table 4, evaluating accuracy on GPQA. Specifically, we assess performance when selecting the top-$k$ agents overall and when choosing $k$ agents at random. In the top-$k$ approach, experts are fixed as the best-performing models for the task, whereas in the random-$k$ strategy, the selected experts vary across instances. Our results demonstrate that skill-based selection is essential. Notably, although selecting the top $k$ experts for a task may seem intuitive, it consistently underperforms compared to Symbolic-MoE’s adaptive instance-level expert selection. Interestingly, top-$5$ selection performs worse than top-$3$ selection, suggesting that a broader selection increases the likelihood of including weaker models, leading to performance degradation. Additionally, the random selection strategy consistently harms performance, showing a $12.86\%$ to $15.61\%$ drop compared to Symbolic-MoE, likely also due to the inclusion of weaker experts. Overall, skill-based instance-level selection consistently outperforms more coarse-grained approaches, with Symbolic-MoE surpassing MoA by $4.92\%$ as also shown in Table 1.
Model MATH-Hard Omni
Debate 83.31 42.93
MoA 86.48 47.36
Self-MoA 88.75 48.75
ReConcile 79.31 42.62
Symbolic-MoE 91.62 52.03
Table 5: Evaluation on MATH-Hard and OmniMATH using profiles from AIME.
Symbolic-MoE generalizes to unseen tasks. Given the constant introduction of new, unseen data in LLM deployments, we further assess whether Symbolic-MoE can generalize to unseen settings using existing model profiles on two additional benchmarks: OmniMATH (Gao et al., 2025) and MATH Hard (Hendrycks et al., 2021b). OmniMATH is an Olympiad-level math benchmark consisting of 4.4k test samples, while MATH Hard is the level 5 (the most challenging) subset of MATH, comprising 1.3k problems. For these benchmarks, we did not create model profiles using their respective validation sets. Instead, we reused the model profiles obtained from AIME. As for the multi-agent baselines, we also use the top-3 models from AIME. Results in Table 5 show that Symbolic-MoE generalizes effectively, even when evaluated on unseen tasks and with model profiles not tailored to the test domain. Notably, the transferred Symbolic-MoE outperforms the best baseline (Self-MoA) by 2.87% on MATH-Hard and 3.28% on OmniMATH.
Aggregator MMLU-Pro GPQA
Random 52.29 48.92
Adaptive 57.12 58.01
Task-specific 63.71 57.78
Table 6: Ablations on different aggregators in our full setting.
Role and selection of the aggregator. Unlike most of our discussion-based multi-agent baselines, Symbolic-MoE collects a single CoT and answer from each expert and combines them via an aggregator. This provides efficiency gains, as shown in Table 2; here, we investigate the role of aggregator in our framework. While experts are selected per instance, the aggregator is chosen per task, as we find that reasoning ability does not necessarily translate to effective aggregation. Table 6 examines the impact of aggregator selection in Symbolic-MoE, comparing three strategies: (1) a randomly chosen aggregator from the model pool (Random), (2) an instance-level aggregator selection based on model profiling (Adaptive), and (3) a task-specific aggregator determined by task-level performance (Task-specific). Evaluated on MMLU-Pro and GPQA, the results indicate that a random aggregator substantially degrades performance, showing that the aggregator plays a crucial role. While the instance-level aggregator improves outcomes on both datasets, the task-specific aggregator outperforms it on MMLU-Pro and performs comparably on GPQA. We further find that the similar performance of instance-specific and task-specific aggregation on GPQA is due to a high degree of overlap in selected aggregators. Overall, this suggests that good reasoners will not always be good aggregators, supporting our task-based selection.
Expert Aggregator GPQA
Random Task-Specific 31.82
Recruited Random 51.52
Recruited Majority Vote 53.54
Recruited Task-Specific 57.78
Table 7: Accuracy on GPQA using different expert selection and aggregation strategies.
Synergy between expert and aggregator selection. As demonstated that the selection of an aggregator plays an important role, we further investigate the synergy between good aggregator selection and good expert selection. In Table 7, we experiment with two expert selection strategies: (1) randomly recruiting $k$ experts without using model profiles, and (2) using our model profiles to recruit experts. We also employ three settings for aggregator selection: (1) based on the task performance as Symbolic-MoE uses, (2) randomly selecting an aggregator, and (3) use majority vote to select the final answer without using any aggregator. Our findings indicate that the combination of strong models with a task-specific aggregator yields the highest performance. When the aggregator is suboptimal, majority voting can serve as a robust alternative. However, when the expert models themselves are weak (chosen randomly), even a strong aggregator cannot compensate for the performance drop.
Keyword LLM MMLU-Pro GPQA
Llama 3.1 8B 64.19 56.62
Gemma 2 9B 64.02 57.01
Qwen 2.5 7B 63.71 57.78
Table 8: Keyword LLM has little effect on the final performance.
Sensitivity to the keyword LLM. We choose Qwen 2.5 7B (Qwen Team, 2024) as the “Keyword LLM” to generate the required skills for each instance during both preprocessing and inference. Here, we investigate the sensitivity of the results to the choice of the Keyword LLM, testing three different models: Qwen 2.5 7B (Qwen Team, 2024), Llama 3.1 8B (AI, 2024), and Gemma 2 9B (Team et al., 2024b). As shown in Table 8, the final performance remains consistent regardless of the chosen model, indicating that the selection of the Keyword LLM has minimal influence on performance.
5 Discussion and Conclusion
A key feature highlighted in Table 1 is the consistency of Symbolic-MoE’s performance. While baseline methods occasionally do well in isolated settings (e.g. MoA on MMLU-Pro, ReConcile on MedMCQA) it is important to highlight that no baseline does well consistently across settings. This means that – without Symbolic-MoE – getting a strong overall result would require evaluating all the baseline methods and choosing the best settings manually. In contrast, Symbolic-MoE achieves high performance without human intervention. By automatically recruiting the experts based on skills needed for each instance, Symbolic-MoE provides a robust recipe that generalizes across domains.
Modularity. Another key advantage of Symbolic-MoE is its modularity. Unlike typical Mixture-of-Experts (MoE) frameworks, which need to be trained end-to-end from scratch in a centralized manner, Symbolic-MoE uses the symbolic output channel of existing models to combine experts. This gradient-free approach enables seamless integration of pre-trained models without updates, allowing them to be trained independently and distributedly. Such delegation enhances domain specialization for each model. Moreover, while standard MoEs have a fixed size determined before training, Symbolic-MoE can adaptively grow and evolve as models are updated. Given the rapid advancements in LLMs, cost-effective and efficient updates are essential – state-of-the-art models are often replaced within months. Symbolic-MoE’s modular and gradient-free design simplifies the incorporation of these updates, requiring only a few calls to obtain a new model’s profile, which can then be easily plugged into this framework. It is also straightforward to increase the number of experts recruited at test time, which is usually fixed in the typical MoE setting.
Connections to Inference-Time Scaling. Like other LLM-discussion frameworks, Symbolic-MoE can be seen as a form of multi-model inference-time scaling. Past works (Snell et al., 2024; Shao et al., 2024) have highlighted the benefits of adding inference-time compute, ranging from simply sampling multiple responses (as in our SC baselines) to more complex strategies such as refinement (Madaan et al., 2023a; Chen et al., 2024b). Symbolic-MoE surpasses several such baselines (Table 1) by adaptively selecting the optimal set of expert models while avoiding the costly discussion process, reducing overhead while improving performance. Moreover, our novel batching method enables Symbolic-MoE to run efficiently on a single GPU and flexibly scale across multiple GPUs.
Conclusion. We introduced Symbolic-MoE, a scalable MoE framework that combines models through their symbolic outputs (i.e., via natural language discussion). Symbolic-MoE infers which skills are needed for a given problem and recruits agents accordingly to engage in a discussion about a given input, guiding the discussion to reach a better consensus. On four diverse reasoning datasets, Symbolic-MoE outperforms standard inference-time scaling methods, debate frameworks and recent mixture-of-agents methods, leading to consistently strong performance across domains without human intervention. Symbolic-MoE’s average performance across heterogeneous tasks is in fact stronger than that of advanced proprietary models such as GPT4o-mini. Moreover, unlike past work that requires multiple models to be loaded on separate GPUs running in parallel, Symbolic-MoE introduces a novel batching strategy that allows us to run on a single GPU in roughly the same amount of time, obtaining the best of both worlds in terms of performance and efficiency.
Limitations
Like other multi-agent discussion methods (Du et al., 2023; Chen et al., 2024c), Symbolic-MoE involves running multiple models, which increases inference cost. This cost can be reduced via distillation: Chen et al. (2024a) distill multi-agent discussions between a fixed set of agents into a single model, showing improvements over distilling from single models. This approach could easily be adapted to distill from a variable set of agents, allowing the student model to benefit from the routing and skill selection process. We leave distilling from Symbolic-MoE to future work.
Symbolic-MoE also relies on skills inferred from a small validation set to set the agent profiles. In our experiments, we ensure fair comparisons to the baselines by choosing models for the baselines according to the same validation set, giving the baselines equal access to the data. Skill inference relies on the Keyword LLM being sufficiently trained on a given domain to infer relevant skills – empirically, we find that this is the case across a variety of domains. Overall, Symbolic-MoE will continue to improve with better skill inference modules, which can easily be swapped in.
Acknowledgements
This work was supported by NSF-CAREER Award 1846185, DARPA ECOLE Program No. HR00112390060, Microsoft Accelerate Foundation Models Research (AFMR) grant program, NSF-AI Engage Institute DRL-2112635, National Institutes of Health (NIH) under other transactions 1OT2OD038045-01, and Cisco and Capital One Faculty Awards. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing official policies, either expressed or implied, of the NIH or other sponsors.
Appendix
Appendix A Illustrations of Variations in Batch Inference
As discussed in Section 3.3, there are mainly three different ways to serve multiple LLMs to solve every instance adaptively. We illustrate these variations in Fig. 3. Our batch inference method allows for the speedups featured in Table 2.
Figure 3: Different approaches to achieving adaptiveness in Symbolic-MoE, which uses different models for each instance. In a naive setup (I), $k$ GPUs must be hosted simultaneously, allowing immediate access to outputs from each model. Another naive setup (II) requires only a single GPU but involves constant loading and offloading of models to obtain outputs from the corresponding model. Our scalable batch inference process (III) strikes a balance between (I) and (II). When models are assigned to problems, we group samples by model and sequentially load the corresponding LLM onto a single GPU to generate outputs efficiently. Moreover, this approach still allows us to parallelize across GPUs if they are available.
Appendix B Baselines Details
We compare against four categories of baselines.
  • Zero-shot single-model methods: This category includes proprietary models such as GPT-4o-mini (OpenAI, 2024), Gemini 1.5 Pro (Team et al., 2024a), and DeepSeek-V3 (DeepSeek-AI et al., 2025b); high-capacity open-source models like Qwen2.5 72B (Qwen et al., 2025) and Llama 3.3 70B (AI@Meta, 2024); and strong distilled 7B models such as QwenR1 (DeepSeek-AI et al., 2025a). For reference, we also report the best task-specific model from our pool for each task, denoted as Task-Best.
  • Advanced single-model baselines with inference-time compute: We evaluate methods that enhance inference-time reasoning, specifically Self-Refine (SR) (Madaan et al., 2023b) and Self-Consistency (SC) (Wang et al., 2023b). To ensure a fair comparison, we set SC’s sample size to 5, aligning with the number of large language model (LLM) calls in Symbolic-MoE, which engages three experts and one aggregator model. [^3] We use an odd number of SC calls to avoid ties. Additionally, for these baselines, we use the best-performing LLM for each task, inferred on the same dev set used for our agent profile creation.
  • Single-model multi-agent baselines: To isolate the impact of Symbolic-MoE’s recruitment strategy, we compare against methods where multiple instances of the same model collaborate. Specifically, we consider Multi-Agent Debate (Debate) (Du et al., 2023) and Self-Mixture-of-Agents (Self-MoA) (Li et al., 2025), both of which rely on iterative, multi-round discussions using a single model. These baselines employ three agents, each using the same task-best model, and conduct two rounds of discussion, resulting in a total of $6$ LLM calls per sample.
  • Multi-model multi-agent baselines: We also evaluate approaches leveraging diverse models in a multi-agent setup. This includes Mixture-of-Agents (MoA) (Wang et al., 2024a) and ReConcile (Chen et al., 2024c), both of which incorporate a fixed set of models in multi-round interactions. To ensure a fair comparison with our approach, particularly in the use of the validation set, we select the top three performing models from the validation set and conduct multi-round interactions. In MoA, agents participate in two rounds of discussion, while agents in ReConcile engage in three rounds, leading to $6$ and $9$ LLM calls per sample, respectively.
Appendix C Implementation Details
We conduct our experiments for Symbolic-MoE and other single-model baselines on a single A6000 GPU with 48 GB of memory, while MoA and ReConcile are executed on 8 A6000 GPUs for parallelization. For the 70B models, we use the original version without quantization and perform inference on 4 A6000 GPUs. All open-source models utilize vLLM (Kwon et al., 2023) for inference. The temperature is set to 0.7 for all methods. The maximum output token length is fixed at $4096$ for all models, except for QwenR1 and LlamaR1, which have a limit of $32768$ since they are trained with longer trajectories and tend to generate longer outputs. All results, except those from proprietary models (due to budget constraints), are averaged over three random seeds. Further details on the model pool, distribution of the expert recruited, and all the prompts we use can be found in Table 9 and Appendix L.
Appendix D Model Pool
We provide the full list of our model pool in Table 9, including their names, sizes, and publicly available checkpoints on Huggingface. Most of the model sizes are 7 to 8 billion.
Model Name Size Huggingface Link
BioLlama 8B ContactDoctor/Bio-Medical-Llama-3-8B
DeepSeekMath 7B deepseek-ai/deepseek-math-7b-instruct
Exaone 7.8B LGAI-EXAONE/EXAONE-3.5-7.8B-Instruct
Gemma2 9B google/gemma-2-9b-it
GLM4 9B THUDM/glm-4-9b
Granite 8B ibm-granite/granite-3.1-8b-instruct
InternLM3 8B internlm/internlm3-8b-instruct
Llama3.1 8B meta-llama/Llama-3.1-8B-Instruct
LlamaR1 8B deepseek-ai/DeepSeek-R1-Distill-Llama-8B
Mathstral 7B mistralai/Mathstral-7B-v0.1
Mistral 12B mistralai/Mistral-Nemo-Instruct-2407
Phi3.5-mini 3.5B microsoft/Phi-3.5-mini-instruct
Qwen2.5 7B Qwen/Qwen2.5-7B-Instruct
Qwen2.5-Coder 7B Qwen/Qwen2.5-Coder-7B-Instruct
Qwen2.5-Math 7B Qwen/Qwen2.5-Math-7B-Instruct
QwenR1 7B deepseek-ai/DeepSeek-R1-Distill-Qwen-7B
Table 9: The models constituting the model pool.
Appendix E Performance on the Validation Set
Table 10 shows the performance of each model on the validation set. We highlight the top-1 and top-3 models in bold font and yellow background, respectively. This information is also used for the baselines we compare against in Table 1.
Model MMLU-Pro AIME GPQA MedMCQA
BioLlama 37.71 0.85 27.31 42.86
DeepSeekMath 32.57 3.32 28.11 35.71
Exaone 52.29 25.99 32.13 56.35
Gemma 53.71 7.73 36.95 64.29
GLM 50.29 7.37 30.92 58.33
Granite 43.43 5.92 34.14 56.15
InternLM 43.14 7.91 36.14 55.56
Llama 46.00 6.78 33.73 66.87
LlamaR1 54.29 51.98 56.22 53.37
Mathstral 34.57 3.11 36.55 52.38
Mistral 45.14 1.41 33.73 46.43
Phi 46.57 1.41 47.79 65.87
Qwen 54.00 13.56 37.35 67.06
QwenCode 46.29 9.89 30.52 50.79
QwenMath 31.71 11.13 28.51 36.90
QwenR1 53.43 57.06 51.41 37.90
Table 10: Comparison of model performance on the validation set. The best model on each task is bolded, and the top 3 models on each task are highlighted in yellow.
Appendix F Performance of Each Model as an Aggregator
Table 11 shows the performance of each model when acting as an aggregator. Note that the best-performing model in Table 10 can be different from the best aggregator model in Table 11, motivating us to choose the aggregator based on this synthetic task described in Section 3.2.2.
Model MMLU-Pro AIME GPQA MedMCQA
BioLlama 37.31 21.47 30.12 42.46
DeepSeekMath 32.57 5.37 21.69 35.71
Exaone 57.43 47.92 35.34 52.58
Gemma 49.71 3.11 31.73 53.37
GLM 52.57 26.27 35.34 51.39
Granite 48.86 36.44 38.96 48.02
InternLM 55.14 16.95 42.57 51.59
Llama 51.14 11.86 40.56 50.60
LlamaR1 59.71 53.67 46.18 49.01
Mathstral 41.71 26.27 35.74 46.43
Mistral 48.00 18.93 33.33 46.43
Phi 27.71 9.04 26.10 25.40
Qwen 56.86 38.14 39.36 53.37
QwenCode 51.14 29.66 38.96 50.79
QwenMath 31.71 5.93 16.06 36.90
QwenR1 58.00 57.63 48.59 45.44
Table 11: Performance of each model when used as an aggregator, on the validation set. The best model on each task is bolded, and is selected as the task-specific aggregator.
Appendix G Distribution of Experts
We present the distribution of recruited experts across different datasets in Fig. 4. As noted in Section 3.2.1, we trim experts with occurrences below 5% to reduce model loading time. In Fig. 4, the top row shows the distribution before trimming, and the bottom row shows the distribution after trimming. The distribution varies significantly across datasets – on more diverse datasets such as MMLU-Pro, the recruited experts are also more varied. In contrast, for AIME and GPQA, which focus more on math and science, the recruited experts are dominated by a few models.
Figure 4: Distribution of the recruited experts across datasets. Top row: the distribution before trimming. Bottom row: the distribution after trimming and re-sampling.
Appendix H Test-time Token Count Analysis
In addition to measuring GPU run time in Table 2, we compare the test-time token count with multi-agent baselines. As shown in Fig. 5, Symbolic-MoE uses fewer tokens than Self-MoA while achieving a significant performance gain. However, compared to MoA and ReConcile, Symbolic-MoE generates more tokens, particularly on GPQA. The primary reason, as illustrated in Fig. 4, is that Symbolic-MoE predominantly recruits LlamaR1 and QwenR1, both of which are trained with long reasoning trajectories, resulting in substantially longer outputs compared to other models. This explains why Symbolic-MoE requires less GPU run time despite producing more tokens: by skipping the expensive multi-round discussions, we eliminate the time spent loading and offloading models. However, the inherent verbosity of the R1 models contributes to the higher token count.
Figure 5: Comparison of the test-time token count used in different methods.
MMLU-Pro GPQA
Acc $\uparrow$ Time $\downarrow$ Acc $\uparrow$ Time $\downarrow$
w/o Trimming 63.94 18.83 55.26 21.78
w/ Trimming 63.71 12.27 57.78 10.85
Table 12: Trimming the low-frequent experts improves both accuracy and efficiency.
Appendix I The Effect of Trimming and Re-sampling
As described in Section 3.3.1, we trim the recruited experts if their occurrence falls below $5\%$ of the total selections. Here, we analyze the impact of this trimming process. Without trimming, the diversity of selected experts increases, but the model loading time also increases. Table 12 presents a quantitative comparison of accuracy and GPU run time using 4 GPUs. As expected, trimming reduces GPU run time across both datasets by minimizing the need to load infrequently used models. Interestingly, we also observe that trimming improves accuracy on GPQA. This improvement may be due to the fact that after trimming, only LlamaR1 and QwenR1 remain as the recruited experts, which are particularly effective on this task.
Appendix J Dataset Statistics and Licenses
We provide the sample sizes and licenses of the datasets used in this work in Table 13. All the datasets are in English, and all datasets are used in a fashion consistent with their intended use.
Validation Size Test Size License
MMLU-Pro (Wang et al., 2024c) 350 2,100 Apache License
AIME (MAA, 2024) 354 30 CC0
GPQA (Rein et al., 2023) 249 198 MIT License
MedMCQA (Pal et al., 2022) 504 4,183 MIT License
Table 13: The statistics and licenses of the datasets we use in this work.
Appendix K Algorithm
We provide the algorithm for our batched inference strategy in Algorithm 1.
Algorithm 1 BatchedInference
1:Test samples $\mathcal{Q}$, Model pool $\mathcal{M}$
2:Inference results for all samples
[…]
4:for $q\in\mathcal{Q}$ do
5:     $E^{(1)}_{q},E^{(2)}_{q},...,E^{(k)}_{q}\leftarrow\textsc{RecruitExperts}(q,\mathcal{M})$triangleright Select $k$ experts per sample (§$\S$3.3.1)
6:     for $e\in E_{q}$ do
7:         expert_sample_map[$e$] $\leftarrow$ expert_sample_map[$e$] $\cup\{q\}$
8:     end for
9:end for
10:
11:$\textit{results}\leftarrow\emptyset$triangleright Results collection
[…]
14:end for
15:return results
Appendix L Prompts
{user_example}
[frametitle=Prompt for the Keyword LLM to Generate Keywords] Question: {question}
What are the core knowledge, subjects or skills needed to solve this problem? List 2-5 keywords separated in comma. Example keywords: psychology, virology, behavioral theory, microbiology, diplomacy, political science, property law, finance, business. Give ONLY the keywords, no other words or explanation.
Follow this format: Keywords: <keyword1>, <keyword2>…
{user_example}
[frametitle=Prompt for Zero-shot Chain-of-Thought Generation (Multiple Choice)] Question: {question}
Provide your step-by-step reasoning first, and then print “The answer is (X)” where X is the answer choice (one capital letter), at the end of your response.
{user_example}
[frametitle=Prompt for Zero-shot Chain-of-Thought Generation (Math)] Question: {question}
[…]
{user_example}
[frametitle=Prompt for the Aggregator (Wang et al., 2024a)] You have been provided with a set of responses from various open-source models to the latest user query. Your task is to synthesize these responses into a single, high-quality response. It is crucial to critically evaluate the information provided in these responses, recognizing that some of it may be biased or incorrect. Your response should not simply replicate the given answers but should offer a refined, accurate, and comprehensive reply to the instruction. Ensure your response is well-structured, coherent, and adheres to the highest standards of accuracy and reliability.
Responses from models:
{model_1_response}
{model_2_response}
{model_3_response}
Question: {question}
Provide your step-by-step reasoning first, and then print “The answer is (X)” where X is the answer choice (one capital letter), at the end of your response.
Appendix M Symbolic-MoE as a Sparse Mixture-of-Expert
In the Sparse Mixture-of-Experts (SMoE) framework (Shazeer et al., 2017a), a trainable router dynamically selects a subset of experts for each input. Formally, given an input $x$, the output of an SMoE layer, $y$ is computed as:
$\begin{split}{y}&=\sum_{i=1}^{k}\mathcal{R}({x})_{i}\cdot f_{i}({x}),\\ \mathcal{R}({x})&=\text{softmax}(\texttt{Top-K}(g(x)),k)\\ \end{split}$ (1)
where $f_{i}(x)$ represents the response of the $i$-th expert, and $\mathcal{R}(x)$ is a trainable router that assigns selection probabilities to each expert based on $g(x)$, typically a small feedforward network (Shazeer et al., 2017b; Riquelme et al., 2021). The Top-K operation retains only the top $k$ experts, setting the probabilities of others to zero after the softmax operation.
However, directly applying SMoE in our framework presents key challenges. Unlike SMoE, our method operates in a symbolic, text-based space and is designed for test-time inference, meaning that we do not rely on a trainable router to learn expert selection, nor do the experts in our method refer to model parameters. Instead, we introduce a skill-based routing mechanism to select relevant experts based on predefined competencies rather than learned gating functions. Formally, our aggregation process can be expressed as:
${\begin{split}y&=A^{*}(\big{\|}_{i=1}^{k}y^{(i)})\\ y^{(i)}&=E^{(i)}(x)~{}\forall~{}i\in\{1,2,...,k\}\\ E^{(i)}&\sim\text{Categorical}(w^{(1)},w^{(2)},...,w^{(n)})~{}\forall i\leq k\\ \end{split}}$ (2)
[…]
This skill-based routing framework retains the core benefits of SMoE while removing the reliance on a trainable gating mechanism. Specifically, the aggregator model $A^{*}$ in Symbolic-MoE plays a role analogous to the weighted sum ($\sum$) operation in SMoE, synthesizing outputs from selected experts. Likewise, the recruited agent $E^{(i)}$ corresponds to the Top-k operation in SMoE, ensuring that only the most relevant and specialized experts contribute to the final output. We inherit the key conceptual benefits of SMoE – dynamic expert selection and response aggregation – while also introducing additional advantages. Symbolic-MoE is gradient-free, eliminating the need for retraining, and is entirely automatic, leveraging a large pool of pre-trained models to deliver a better performance.